f(x)=2x^2-x+1
use the discrimant to determine whether the function has complex zeros

Answer by jim_thompson5910(35256)

D = b^2 - 4ac


D = (-1)^2 - 4(2)(1)


D = 1 - 4(2)(1)


D = 1 - 8


D = -7


Since the discriminant is negative, f(x) will have two complex zeros.